CRY+VIS Interface Meeting
Date: 2022/6/10 13:00-13:25
Participants: Takahashi, Yokozawa, Uchiyama, Ushiba, Ikeda, Aso, Sato, Miyoki, Akutsu, Kimura, Hirata
- EX: Height adjustment of ETMX. Centering of QPD in TMSX.
- IY: Leak test in the next week.
- EY: Offload of GAS filters.
- Type-B: Replacement of H2 geophone in SR2. Stand-alone SM driver for SR3.
- Type-Bp: FC on PR3. Installation of traverser limit SW in PR3. For PR2 in 6/14-15. MSB to PR3 in 6/16-17.
- Type-C: Alignment of OMC. Recovery of OMMT2 suspension. Adjustment of H3 OSEM in OMMT1 after the final alignment (if possible).
- Replacement of flip cables: EXT and BF temp.
- TM height setting: Following the cooling.
- Oplev: Investigating the relation between laser and temperature.
- Commissioning in air:
- Commissioning in vacuum: GR in Y-arm.
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: Fan was attached to IYV pump. Test in the next week.
- Heater control: Pending.
- TMS: QPD in EX.
Date: 2022/6/10 14:00-14:15
Participants: Takahashi, Uchiyama, Ushiba, Fabian, Tanaka, Aso
- Adjusted the height of ETMX.
- Offloaded the GAS filters in EY and IY.
In SR2 we identified that geophone 2 had a problem, so we replaced it with a spare one. Takahashi-san gave me the information about calibration factors, but I still have to take a look at it.
- We installed new cable clamps and limit switches in PR3.
- We locked the suspension for First Contact application.
- Recovered the IM touching the magnets in OMMT2 suspension.
- Performed the health check of OMMT1, OMMT2, and OSTM after the OMC alignment. It is necessary to adjust the H3 OSEM in OMMT1.
- Type-B: Stand-alone SM driver for SR3, Geophone work in SR2.
- Type-Bp: MSB on PR3. Traverser limit SW in PR2.
(Date in Kamioka)
The week from 6/13
- Takahashi: 6/14-17
- Sato: 6/14-17
- Hirata: 6/14-17
The week from 6/20